Given the aggregate consumption function C=0.9y+100 find the mpc and apc and find elasticity of consumption with respect to income show that it equals mpc and apc
1
Expert's answer
2020-10-02T06:54:17-0400

Solution

Given function

C=0.9y +100

Slop of this curve will give mpc

So

MPC=0.9


APC of a consumption function is given as

APC=C/Y

SO at y=10 c=109 from consumption function so

APC=109/10=10.9

IF y=20 then c=118 then

APC=5.9

IF y=50 then c=145 so

APC=145/50=2.9

Elasticity of consumption with respect to income can be calculated as

% change in consumption="\frac{(118-109)}{109}\times100=8.25%" %

%change in price="\frac{(20-10)}{10}\times100=100" %

Elasticity of consumption with respect to income ="\frac{8.25}{100}=0.08"